OpenStudy (anonymous):

converge or diverge {sum_{1}^{infty} 3^{-2} e ^{k}}

OpenStudy (anonymous):

explain

OpenStudy (experimentx):

\[ {\sum_{k=1}^{\infty} 3^{-2} e ^{k}} \] Looks much better, use similar test as before ... it's quite easier that the last one.

OpenStudy (anonymous):

can i pull the constant out

OpenStudy (experimentx):

yes

OpenStudy (anonymous):

series diverges because limit approaches infinity

OpenStudy (turingtest):

this seems too obvious, makes me suspicious something is missing

OpenStudy (anonymous):

i thought there was more to it as well but idk

OpenStudy (experimentx):

looks like you are missing some part of question.

OpenStudy (anonymous):

nope thats the question

OpenStudy (turingtest):

it's just crazy obvious with a positive exponent that\[\frac19\sum_{k=1}^{\infty}e^k\]diverges I would think they would maybe ask you about\[\frac19\sum_{k=1}^{\infty}e^{-k}\]

OpenStudy (experimentx):

it clearly looks divergent.

OpenStudy (anonymous):

thanks guys